Circular No. 7567

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2000 Y6 AND C/2000 Y7 (SOHO) Further to IAUC 7565, D. Hammer reports his measurements for two

comets (initial observations given below) that appear to be two components of an earlier single comet. C/2000 Y6 and C/2000 Y7 were found by M. Meyer and by S. Hoenig, respectively, in C2 coronagraph data on SOHO website images. D. Biesecker provides V magnitudes for C/2000 Y6: Dec. 20.463 UT, 7.8; 20.504, 7.8; 20.580, 7.5; 20.588, 7.6; 20.604, 8.0; 20.646, 8.3. The reduced observations and parabolic orbital elements (T = 2000 Dec. 20.85 TT, q = 0.025 AU, ω = 88^{o}–89^{o}, Ω = 229^{o}, i = 87^{o}–89^{o}) by B. G. Marsden, together with a search ephemeris for groundbased observers, are given on MPEC 2001-B08.

Comet  2000 UT  α_{2000}  δ_{2000}
C/2000 Y6  Dec. 20.454  17^{h}48.^{m}7  −23^{o}20′
C/2000 Y7  20.463  17 48.8  −23 19

SUNGRAZING COMETS Hammer also reports his measurements for seven additional Kreutz

sungrazing comets, all found in C2 data (except C/2001 A3, in C3 data) at the SOHO website by T. Scarmato (C/2000 J6), X. Leprette (C/2000 K7), M. Oates (C/2000 K8, C/2000 Y9), and M. Boschat (remaining comets). The reduced observations and orbital elements by Marsden appear on the MPECs cited below.

Comet  2000 UT  α_{2000}  δ_{2000}  MPEC
C/2000 J6 May 11.410  3^{h}18.^{m}1  +16^{o}37′  2001-B11
C/2000 K7  19.160  3 47.4  +18 17  2001-B11
C/2000 K8  29.788  4 28.5  +20 07  2001-B11
C/2000 Y5 Dec. 19.736  17 58.8  −25 08  2001-B08
C/2000 Y8  23.788  18 17.7  −24 52  2001-B08
C/2000 Y9  25.121  18 23.6  −24 53  2001-B08
Comet  2001 UT  α_{2000}  δ_{2000}  MPEC
C/2001 A3 Jan. 7.971  19^{h}32.^{m}6  −24^{o}10′  2001-B11
2001 January 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green